# Posterior Ramifications of Prior Dependence Structures

### v2

@@ 1.4 Contributions
This paper provides general guidance for prior dependence specification in multivariate settings. These recommendations are topical given that recent advances in copula-based priors allow for the incorporation of unprecedented flexibility into the prior dependence structure. While this additional flexibility could give rise to priors that more accurately characterize real-life phenomena, practitioners may be subject to choice overload when deciding between the vast number of potential dependence structures. We argue that this additional flexibility is only useful in certain contexts. Our recommendations help practitioners discard prior dependence structures that do not align with their objectives for posterior analysis – simplifying the prior specification process. These recommendations also clarify how the prior dependence structure can impact the posterior distribution. Our recommendations are illustrated using several models for which copula-based priors have been proposed. Unlike situations that require the specification of a community of prior distributions, we restrict our discussion to the case where a single prior is elicited.
The remainder of this article is structured as follows. In Section 2, we define general conditions under which the prior dependence structure is incompatible with that induced by the likelihood function and unable to be retained by the posterior distribution as data are observed. We formally prove this large-sample result and demonstrate that priors elicited using the methods proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) are generally incompatible with standard likelihood functions. We discuss small-sample scenarios where this additional flexibility in the prior dependence structure is nevertheless useful in Section 3. In Section 4, we prove asymptotic results about the impact of the prior dependence structure on the convergence of the posterior mode, which we contrast with previous work on copula-based priors (Michimae and Emura, 2022). We then conduct numerical studies with both small and large sample sizes to contextualize these theoretical results. These simulations prompt further recommendations regarding choosing a prior dependence structure for posterior analysis. We provide concluding remarks and a discussion of extensions to this work in Section 5.
@@ 2.1 Background
In this section, we examine situations where the prior dependence structure cannot carry over into the posterior distribution as data are collected. It is unrealistic to expect the prior dependence structure for $\boldsymbol{\theta}$ to be perfectly specified. However, we should be mindful of whether the prior dependence structure can be retained a posteriori before using the posterior of $\boldsymbol{\theta}$ to draw conclusions about its dependence structure. Otherwise, we may draw different conclusions about the dependence structure given small and large samples generated from the same data generation process. Drawing such conclusions in an uninformed way presents practical issues and complicates the consideration of the posterior of a function of several parameters in $\boldsymbol{\theta}$.
The concept of chronic rejection (Libby and Pober, 2001; Vos et al., 2011) frames this section’s main result. The term chronic rejection describes the process in which a transplanted organ is rejected by the recipient’s immune system over a long period of time. The recipient’s persistent immune response against the transplanted organ causes gradual damage. Similar to screening an organ donor, we argue that one should consider whether the prior dependence structure for $\boldsymbol{\theta}$ is compatible with that induced by the likelihood function. Such incompatibilities imply that the prior dependence structure cannot be retained a posteriori – and the prior dependence structure is hence a chronically rejected one. We emphasize that the term rejection as used in this paper does not imply the formal rejection of a statistical hypothesis test. In Section 2.2, we define the notion of chronically rejected prior dependence structures. We illustrate this definition using simulation in Section 2.3. Section 3 discusses situations where using chronically rejected prior dependence structures may be sensible, so practitioners should consider their posterior objectives before choosing a prior dependence structure.
@@ 2.2 Chronically Rejected Prior Dependence Structures
Here, we define general conditions under which prior dependence structures cannot be retained by the posterior with enough data. These sufficient conditions can be readily verified prior to observing data. When planning posterior analyses, practitioners can discard prior dependence structures that satisfy these conditions for certain posterior objectives. This simplification reduces the number of potential dependence structures and hence streamlines the prior elicitation process.
We consider the limiting behaviour of the posterior distribution for $\boldsymbol{\theta}$. Because we consider this behaviour under various data generation processes, the data are random variables. Data from a random sample of size $n$ are represented by $\boldsymbol{Y}^{{}_{(n)}}$. Realizations of these samples are denoted by $\boldsymbol{y}^{{}_{(n)}}$. Much of the work on limiting posteriors (see e.g., Ghosal et al. (1995), Le Cam and Yang (2000), Gao et al. (2020), or Schillings et al. (2020)) appeals to the Bernstein-von Mises theorem (van der Vaart, 1998). We let $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ be the statistical model corresponding to the likelihood function in (1). When data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently and identically from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$, the Bernstein-von Mises theorem dictates that the posterior for $\boldsymbol{\theta}$ converges to the $\mathcal{N}(\boldsymbol{\theta}_{0},\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n)$ distribution in the limit of infinite data, where $\mathcal{I}(\cdot)$ is the Fisher information.
@@
Under the conditions in Theorem 1, the posterior copula for $\boldsymbol{\theta}$ is approximately Gaussian for large samples. However, we may not collect nearly enough data for this copula to be Gaussian in practice. Because it is unrealistic to expect the prior dependence structure of $\boldsymbol{\theta}$ to be perfectly specified, we consider a partial characterization of this prior (and posterior) dependence structure using D-vines in Corollary 1. Even if the copula in (4) is not specified using a D-vine, such a structure can be induced. D-vines are often specified via the set of Kendall’s $\tau$ (Kendall, 1938) values for each of the D-vine’s bivariate copulas (Kurowicka and Cooke, 2005). Kendall’s $\tau$ measures rank correlation in terms of how similar the orderings of bivariate data are when ranked by each quantity. We let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ characterize the prior dependence structure on $\boldsymbol{\theta}$.
@@
Under the conditions for Theorem 1, let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ describe prior dependence on $\boldsymbol{\theta}$. Suppose no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ is such that $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure $\{\tau(\boldsymbol{\theta})_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ such that $\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}=\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}$ for all $\boldsymbol{e}\in E(\mathcal{V})$. Then, the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{Y}^{{}_{(n)}}$ cannot retain the magnitude and direction of prior dependence as $n\rightarrow\infty$.
@@
Corollary 1 follows directly from Theorem 1. We suppose there is no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$. It follows that the magnitude and direction of the dependence structure for $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ cannot be characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ for sufficiently large $n$. We emphasize that considering dependence structures via Kendall’s $\tau$ on the D-vine structure of the copula does not fully specify the dependence structure. This allows for more flexibility in the choices for the copula families; it also facilitates the consideration of prior dependence for subvectors of $\boldsymbol{\theta}$, which is useful because it may require too much cognitive complexity to assess the full prior dependence structure. However, two bivariate copulas may have the same Kendall’s $\tau$ but different properties in terms of symmetry and tail dependence. We focus on the magnitude and direction of dependence to present generally applicable guidance for prior specification.
@@
Corollary 1 therefore provides a straightforward result that can be used to discard potential dependence structures for $\boldsymbol{\theta}$ that cannot be retained as data $\boldsymbol{y}^{{}_{(n)}}$ are observed. We refer to prior dependence structures that satisfy the conditions for Corollary 1 as chronically rejected dependence structures. The conditions for Corollary 1 are sufficient in that there is no value for $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that these dependence structures will be retained as data are generated via (5). However, it is not guaranteed that the prior dependence structure will be retained when the conditions for Corollary 1 are not satisfied. That is, the true value of value $\boldsymbol{\theta}_{0}$ might not be one of those that prevent the conditions for Corollary 1 from being satisfied. Thus, a prior dependence structure that is not retained is not necessarily a chronically rejected one. In this paper, the notion of chronic rejection defines a specific class of prior dependence structures, whereas retention of the prior dependence structure is viewed more generally as a posterior outcome.
@@ 2.3 Illustration with Copula-Based Priors for the Multinomial Model
We now apply Corollary 1 with the multinomial model. In Appendix A of the supplementary materials, we show that the inverse Fisher information matrix for the multinomial model parameterized in terms of $\boldsymbol{\theta}=(Z_{1},Z_{2},\dots,Z_{w-1})$ from (2) is diagonal for all possible $(Z_{1},Z_{2},\dots,Z_{w-1})\in(0,1)^{w-1}$. If $p(\boldsymbol{\theta})$ incorporates any positive or negative dependence between the conditional multinomial probabilities, this dependence structure is therefore a chronically rejected one.
For $w=3$ categories, we illustrate this phenomenon when combining such priors with the standard multinomial likelihood, which is parameterized by conditional probabilities $\boldsymbol{\theta}=(Z_{1},Z_{2})$. We specify $p(\boldsymbol{\theta})$ as in (4) where the marginal prior for $Z_{1}$ is $\text{BETA}(20,40)$, the marginal prior for $Z_{2}$ is $\text{BETA}(30,30)$, and these marginal priors are joined using a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. This prior distribution conveys that we expect the multinomial probabilities to be roughly equal for all three categories. This prior specification could be facilitated using either of the methods by Elfadaly and Garthwaite (2017) or Wilson (2018). We note that the prior copula gives rise to a value for Kendall’s $\tau$ of $0.7132\text{sin}^{-1}(-0.9)/\pi=-$ 0.713.
@@
This table shows that the posterior cannot retain the negative dependence structure present in the prior as data are observed from the prior predictive distribution. Even though the analysis prior $p(\boldsymbol{\theta})$ coincides with that used to define the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$, the multinomial likelihood function cannot accommodate this dependence between $Z_{1}$ and $Z_{2}$. As the sample size increases, we observe that the impact of the likelihood function on the posterior overwhelms that of the prior, and the prior dependence structure will eventually not be retained when the conditions for Corollary 1 hold.
The copula-based priors proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) define valid posteriors for $\boldsymbol{\theta}$ when combined with the multinomial likelihood, but Table 1 corroborates that we would draw vastly different conclusions about the posterior dependence structure of $\boldsymbol{\theta}$ for small and large samples. Elfadaly and Garthwaite (2017) provided R code to combine their priors with the multinomial likelihood function, and Wilson et al. (2021) stated that these priors could be used for Bayesian analysis. However, neither contribution clearly disclosed that these more flexible prior dependence structures cannot be retained as multinomial data are collected. Given a candidate prior dependence structure, Corollary 1 helps assess whether this dependence structure for $\boldsymbol{\theta}$ can be retained a posteriori. We elaborate on how to determine whether the inability to retain the prior dependence structure presents practical issues in Section 3.

### v3

@@ 1.4 Contributions
This paper provides general guidance for prior dependence specification in multivariate settings. These recommendations are topical given that recent advances in copula-based priors allow for the incorporation of unprecedented flexibility into the prior dependence structure. This additional flexibility could give rise to priors that more accurately characterize real-life phenomena, but we argue that this flexibility is only useful in certain contexts. As such, statisticians should be equipped to help practitioners cull the vast number of potential prior dependence structures and select one that aligns with their objectives for posterior analysis. By exploring how the prior dependence structure can impact the posterior distribution, our recommendations clarify when it is useful to elicit complicated prior dependence structures and draw inferences regarding posterior dependence. These recommendations are illustrated using several models for which copula-based priors have been proposed. Unlike situations that specify a community of prior distributions, we restrict our discussion to the case where a single prior is elicited.
The remainder of this article is structured as follows. In Section 2, we define general conditions under which the prior dependence structure is incompatible with that induced by the likelihood function and unable to be retained by the posterior distribution as data are observed. This result may not surprise Bayesian statisticians. Yet, we argue that it is important to explicate this finding clearly – both for the sake of practitioners and when teaching students who will eventually guide practitioners as statistics experts. In Section 3, we emphasize small-sample scenarios in which flexible prior dependence structures can be useful even if they are not retained asymptotically. We also discuss how the inability to retain prior dependence might be used to diagnose philosophical issues with the objectives for posterior analysis. In Section 4, we prove asymptotic results about the impact of the prior dependence structure on the convergence of the posterior mode, which we contrast with previous work on copula-based priors (Michimae and Emura, 2022). We then conduct numerical studies with both small and large sample sizes to contextualize these theoretical results and prompt further recommendations for choosing prior dependence structures. We provide concluding remarks and a discussion of extensions to this work in Section 5.
@@ 2.1 Background
In this section, we examine situations where the prior dependence structure cannot carry over into the posterior distribution as data are collected. It is unrealistic to expect the prior dependence structure for $\boldsymbol{\theta}$ to be perfectly specified. However, we should be mindful of whether the prior dependence structure can be retained a posteriori before using the posterior of $\boldsymbol{\theta}$ to draw conclusions based on its dependence structure. This awareness primes statisticians for questions from practitioners about why we may draw different conclusions given small and large samples generated from the same data generation process. The inability to retain prior dependence might also allude to misalignment between the complexity of the likelihood function and the objectives for posterior analysis.
The concept of chronic rejection (Libby and Pober, 2001; Vos et al., 2011) frames this section’s main result. The term chronic rejection describes the process in which a transplanted organ is rejected by the recipient’s immune system over a long period of time. The recipient’s persistent immune response against the transplanted organ causes gradual damage. Similar to screening an organ donor, we argue that one should consider whether the prior dependence structure for $\boldsymbol{\theta}$ is compatible with that induced by the likelihood function. Such incompatibilities imply that the prior dependence structure cannot be retained a posteriori – and the prior dependence structure is hence a chronically rejected one. We emphasize that the term rejection as used in this paper does not imply the formal rejection of a statistical hypothesis test. In Section 2.2, we define the notion of chronically rejected prior dependence structures. We illustrate this definition with priors elicited using methods from Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) and Wilson (Wilson, 2018) in Section 2.3. To help inform conversations between statisticians and practitioners, Section 3 discusses whether the use of chronically rejected prior dependence structures aligns with various posterior objectives.
@@ 2.2 Chronically Rejected Prior Dependence Structures
Here, we define general conditions under which prior dependence structures cannot be retained by the posterior with enough data. These sufficient conditions can be readily verified prior to observing data to assess the alignment of the prior dependence structure with posterior objectives. We consider the limiting behaviour of the posterior distribution for $\boldsymbol{\theta}$. Because we consider this behaviour under various data generation processes, the data are random variables. Data from a random sample of size $n$ are represented by $\boldsymbol{Y}^{{}_{(n)}}$. Realizations of these samples are denoted by $\boldsymbol{y}^{{}_{(n)}}$. Much of the work on limiting posteriors (see e.g., (Ghosal et al., 1995), (Le Cam and Yang, 2000), (Gao et al., 2020), or (Schillings et al., 2020)) appeals to the Bernstein-von Mises theorem (van der Vaart, 1998). We let $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ be the statistical model corresponding to the likelihood function in (1). When data $\boldsymbol{Y}^{{}_{(n)}}$ are generated independently and identically from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$, the Bernstein-von Mises theorem dictates that the posterior for $\boldsymbol{\theta}$ converges to the $\mathcal{N}(\boldsymbol{\theta}_{0},\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}/n)$ distribution in the limit of infinite data, where $\mathcal{I}(\cdot)$ is the Fisher information.
@@
Under the conditions in Theorem 1, the posterior copula for $\boldsymbol{\theta}$ generally becomes more limited in its flexibility to characterize dependence. To see this, notice that the posterior copula for $\boldsymbol{\theta}$ is approximately Gaussian for large samples. Moreover, the magnitude and direction of the dependence in this approximately Gaussian copula is constrained by the likelihood function through the matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ and possible values for $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}^{*}$. Yet, we may not collect nearly enough data for this copula to be Gaussian in practice.
Because it is unrealistic to expect the prior dependence structure of $\boldsymbol{\theta}$ to be perfectly specified, we consider a partial characterization of the dependence structure based on Kendall’s $\tau$ values for D-vines in Corollary 1. Even if the copula in (4) is not specified using a D-vine, such a structure can be induced. We let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ defined in Section 1.3 characterize the prior dependence structure on $\boldsymbol{\theta}$. The matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ and standard results about conditional normality prompt limiting Kendall’s $\tau$ values for each pair of variables $e_{1}$ and $e_{2}$ in $\boldsymbol{\theta}$ given the variables in their conditioning set $D_{e}$: $\{\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$.
@@
Under the conditions for Theorem 1, let $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ describe prior dependence on $\boldsymbol{\theta}$. Suppose no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ is such that $C^{Ga}_{\boldsymbol{R}}(\cdot)$ corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure $\{\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ such that $\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}=\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}$ for all $\boldsymbol{e}\in E(\mathcal{V})$. Then, the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{Y}^{{}_{(n)}}$ cannot retain the magnitude and direction of prior dependence as $n\rightarrow\infty$.
@@
Corollary 1 follows directly from Theorem 1. We suppose there is no $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that the Gaussian copula corresponding to the covariance matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ induces a dependence structure characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$. It follows that the magnitude and direction of the dependence structure for $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ cannot be characterized by $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$ for sufficiently large $n$. We emphasize that considering dependence structures via Kendall’s $\tau$ on the D-vine structure of the copula does not fully specify the dependence structure. This allows for more flexibility in the choices for the copula families; it also facilitates the consideration of prior dependence for subvectors of $\boldsymbol{\theta}$, which is useful because it may require too much cognitive complexity to assess the full prior dependence structure. However, two bivariate copulas may have the same Kendall’s $\tau$ but different properties in terms of symmetry and tail dependence. We focus on the magnitude and direction of dependence to present generally applicable guidance for prior specification in Section 3.
@@
Corollary 1 facilitates the straightforward identification of dependence structures for $\boldsymbol{\theta}$ that cannot be retained as data $\boldsymbol{y}^{{}_{(n)}}$ are observed. This identification can diagnose potential disagreement between the degree of flexibility in the prior dependence structure and that induced by the likelihood function. We refer to prior dependence structures that satisfy the conditions for Corollary 1 as chronically rejected dependence structures. The conditions for Corollary 1 are sufficient in that there is no value for $\boldsymbol{\theta}_{0}$ with $p_{D}(\boldsymbol{\theta}_{0})>0$ such that these dependence structures will be retained as data are generated via (5). However, it is not guaranteed that the prior dependence structure will be retained when the conditions for Corollary 1 are not satisfied. That is, the true value of $\boldsymbol{\theta}_{0}$ might not be one of those that prevent the conditions for Corollary 1 from being satisfied. Thus, a prior dependence structure that is not retained is not necessarily a chronically rejected one. In this paper, the notion of chronic rejection defines a specific class of prior dependence structures, whereas retention of the prior dependence structure is viewed more generally as a posterior outcome.
@@ 2.3 Illustration with Copula-Based Priors for the Multinomial Model
We now apply Corollary 1 with the multinomial model. In Appendix A of the supplementary materials, we show that the inverse Fisher information matrix for the multinomial model parameterized in terms of $\boldsymbol{\theta}=(Z_{1},Z_{2},\dots,Z_{w-1})$ from (2) is diagonal for all possible $(Z_{1},Z_{2},\dots,Z_{w-1})\in\boldsymbol{\Theta}^{*}=(0,1)^{w-1}$. That is, $\tau(\boldsymbol{\theta}_{0})_{e_{1},e_{2}|D_{\boldsymbol{e}}}=0$ for all $\boldsymbol{e}\in E(\mathcal{V})$ and $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}^{*}$. For distributions in the exponential family (Lehmann and Casella, 1998), conditions 1, 2, and 3 for the Bernstein-von Mises theorem are satisfied as long as the prior $p(\boldsymbol{\theta})$ has support over $\boldsymbol{\Theta}^{*}$. If $p(\boldsymbol{\theta})$ incorporates any positive or negative dependence between the conditional multinomial probabilities, then $\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\neq 0$ for some $\boldsymbol{e}\in E(\mathcal{V})$. By Corollary 1, all such dependence structures are therefore chronically rejected ones.
For $w=3$ categories, we illustrate this phenomenon when combining such priors with the standard multinomial likelihood, which is parameterized by conditional probabilities $\boldsymbol{\theta}=(Z_{1},Z_{2})$. We specify $p(\boldsymbol{\theta})$ as in (4) where the marginal prior for $Z_{1}$ is $\text{BETA}(20,40)$, the marginal prior for $Z_{2}$ is $\text{BETA}(30,30)$, and these marginal priors are joined using a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. This prior distribution conveys that we expect the multinomial probabilities to be roughly equal for all three categories. This prior specification could be facilitated using either of the methods by Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) or Wilson (Wilson, 2018). We note that the prior copula gives rise to a value for Kendall’s $\tau$ of $0.7132\text{sin}^{-1}(-0.9)/\pi=-$ 0.713.
@@
This table shows that the posterior cannot retain the negative dependence structure present in the prior as data are observed from the prior predictive distribution. Even though the analysis prior $p(\boldsymbol{\theta})$ coincides with that used to define the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$, the multinomial likelihood function cannot accommodate this dependence between $Z_{1}$ and $Z_{2}$. Corollary 1 identifies this disagreement between the level of flexibility in the prior dependence structure and that induced by the simple multinomial likelihood function before observing data.
@@
The copula-based priors proposed by Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) and Wilson (Wilson, 2018) define valid posteriors for $\boldsymbol{\theta}$ when combined with the multinomial likelihood, but Table 1 corroborates that we would draw vastly different conclusions about the posterior dependence structure of $\boldsymbol{\theta}$ for small and large samples. While Bayesian statisticians may anticipate such results, they may be unexpected by practitioners whom these results materially impact when deciding between prior dependence structures and how much effort to invest in the elicitation process. Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) provided R code to combine their priors with the multinomial likelihood function, and Wilson et al. (Wilson et al., 2021) stated that these priors could be used for Bayesian analysis. However, neither contribution clearly disclosed that these more flexible prior dependence structures cannot be retained as multinomial data are collected. This example segues into a broader discussion about when the inability to retain the prior dependence structure presents or reveals practical issues with posterior analysis in Section 3.